Determine whether the following Differential equation is seperable, Linear, Exact, homogeneous, and Bernoulli. Next find its solution by respective method.


8 cos^2 ydx + csc^2 xdy=0

1
Expert's answer
2021-11-01T18:53:46-0400

Rewrite in the form of a first order separable ODE


"\dfrac{1}{\sin^2x}dy+8\cos^2ydx=0"

"\dfrac{1}{\cos^2y}dy=-8\sin^2xdx"

Integrate


"\int\dfrac{1}{\cos^2y}dy=-\int8\sin^2xdx"

"\tan y=-4\int(1-\cos(2x)dx"

"\tan y=-4x+2\sin(2x)+C"

"y=\tan^{-1}(-4x+2\sin(2x)+C)+\pi n, n\in\Z"